# Quantum Neural Networks under Depolarization Noise: Exploring White-Box Attacks and Defenses

### 29 Nov 2023 v1

@@ 3.2 Model Architecture
Table 1: Hyperparameters used for the various networks for the different datasets. Note that for the circuit estimate one needs to also include the amplitude encoding procedure by Mottonen et al. [22].
Hyper-Parameter Iris MNIST2 MNIST4 MNIST10 learning rate 0.05 0.005 0.005 0.005 batch size 30 512 512 128 Nr. epochs 100 30 30 30 Nr layers 2 40 40 40 Nr. measurements 1 1 2 4 Nr. Parameters 13 961 961 961 Circuit Depth 7 196 196 196 Single Qubit Gates 12 960 960 960 Two Qubit Gates 4 320 320 320
@@
Additionally, due to the noisy architecture and limited availability of quantum devices, we have simulated the quantum device using statevector simulation. Specifically, we utilized the auto-grad feature of pytorch and the pennylane library to formulate the ansatz  [^5] Pennylane 0.29.1 and pytorch 2.0.0.
@@ 4 Experiments
With our experiments, we try to answer the three questions raised above. At first, we want to see the influence of depolarization noise on a multi-class classifier. Secondly, our interest shifts towards whether one can defend our QNNs with other adversarial training methods, specifically randomized smoothing and adversarial training using PGD. Finally, we want to elaborate on whether we can chain adversarial training with depolarization noise as a blend of two adversarial defense mechanisms.
@@
Dataset Train Param. Acc. Certified Distance: $\tau_{D}$ (×$\times 10^{-2}$)
@@ 4.1 Training of the Models
Throughout this study, we trained four types of QNNs: one for the Iris dataset, one for the MNIST-2, one for the MNIST-4, and one for the MNIST-10 dataset. We train all models using Adam [14] as an optimizer by varying the batch size, learning rate, and number of epochs. We will refer to those models as vanilla models. We used categorical-cross entropy implementation from pytroch as a loss function for Iris, MNIST2 and MNIST4. For MNIST10, we found that simply using negative log likelihood ($l(y,\hat{y})=-\sum_{i}y_{i}\log(\hat{y}_{i})$) as a loss further increased the accuracy, which is why this loss function was used. Concerning adversarial training, we train the models using the PGD attack with $L_{\infty}$-norm for the Iris dataset with $T=50$ steps. Due to performance issues, we train the MNIST-2, MNIST-4, and MNIST-10 models with $T=3$ steps. We additionally employ randomized smoothing during training; therefore, we perturb the original image with noise from a uniform distribution and use this as a training sample. We report the variance of this training as $\sigma$, and the mean is set to zero. We employ early stopping for both randomized smoothing and adversarial training. Therefore, we use the loss value, at which we considered the models to be converged. For Iris this value was set to $0.6$. For MNIST2, we set this value to $0.45$, for MNIST4 to $1.15$ and for MNIST10 to $1.7$. Table 2 reports the test accuracy, combined with the mean certifiable distance $\tau_{D}$ we derived from the test-set.
4.2 Effect of Depolarization Noise on a Multi-Class Classifier
We compare the vanilla networks of the Iris, MNIST2, MNIST4, and MNIST10 datasets for depolarization noises $p$ in the range $[0.1,0.5,0.9]$. We want to point out that we have included a depolarizing noise of $0.3$ for MNIST2. Our results are shown in Figure 4.
Overall, the two binary classification networks, MNIST2 and Iris, perform well in terms of plain accuracy on the test-set. Depolarization noise might help here in the sense that it improves the classifier for larger epsilon values, which we can observe for $p=0.5$ and for $p=0.9$ for the Iris dataset. Similar behavior can also be seen in Du et al. [6, Apdx. H], where the adversarial accuracy was evaluated under a $L_{2}$-Norm. For MNIST2, we did not find any significant improvement in terms of adversarial accuracy. Similarly, for MNIST4 and MNIST10, we can observe that depolarization noise will not increase the robustness. We account this fact towards two things.
The first issue concerns the limited capacity of the network to learn the data distribution. This observation is supported by Table 2. We can see that the accuracy further decreases with an increase in the number of classes. Additionally, when looking at the certified bounds, we can first see that the minimum of all computed bounds is typically negative, and the maximum trace distance is typically smaller than for the binary classification. The negative bound can be attributed to our inability to achieve a $100\%$ accuracy on the test set.
@@ 4.2 Effect of Depolarization Noise on a Multi-Class Classifier
(b) MNIST2
(c) MNIST4
(d) MNIST10
Figure 4: Robustness accuracy for the various vanilla classifiers for Iris, MNIST2, MNIST4 and MNIST10.
Secondly, we have measured two qubits to obtain the classes for MNIST4 and four qubits for MNIST10. Another attribution of the bad performance could be, that depolarization noise will this way automatically increase the probability of the follow-up class since we are assigning two classes to one qubit. Nevertheless, this assumption requires a further experimental and analytical evaluation, which is subject for future work.
@@ 4.3 Effect of Adversarial Training
Our results, regarding the adversarial training of the QNNs can be found in Figure 5. Overall, we can see, that neither of the techniques provides a reasonable estimate of enhancing the robustness of the QNN. Additionally, we can see a similar behavior as shown for the depolarizing noise, specifically that the networks tend to get less robust for larger classes. This development shows a parallel trend towards the fact that the testing accuracy drops around $\approx 10\%$ from MNIST2 to MNIST4 and around $\approx 10\%$ from MNIST4 to MNIST10. Interestingly, randomized smoothing and adversarial training tend not to worsen the robustness accuracy.
@@
(c) MNIST2: PGD
@@
(e) MNIST4: PGD
(f) MNIST4: RS
(g) MNIST10: PGD
(h) MNIST10: RS
Figure 5: Comparsion of the two methods randomized smoothing (RS) and adversarial training (PGD) for a set of parameters on the QNN. Adversarial training was conducted with PGD($L_{\infty}$, $T=50$) for the Iris classifier, PGD($L_{\infty}$, $T=3$) for all MNIST classifiers
4.4 Effect of Depolarization Noise to Adversarial Training
Since adversarial training and the effect of depolarization noise tend to behave similarly, we want to explore if adding depolarization noise on top of an adversarial training procedure can improve the overall robustness of the network. Therefore, we look closer at the MNIST2 dataset and enhance the adversarially trained QNN circuits with depolarization noise while attacking them. Our results can be observed in Figure 6. Overall, we find that adding depolarization noise on top of adversarial training does not show the ability to provide improvements for the MNIST2 Dataset. We can see a similar picture for other datasets when looking at the certified trace distances on the test set in Table 2.
@@ 4.4 Effect of Depolarization Noise to Adversarial Training
5 Related Work
Since quantum neural networks are prune to adversarial attacks [18, 19], over recent years, multiple authors have tried to use quantum computing techniques to enhance adversarial robustness for both classical and quantum neural networks. Consequently, Sahdev and Kumar [26] have utilized a quantum phase estimation to provide a quadratic speedup for randomized smoothing. Nevertheless, since quantum phase estimation is not feasible to be executed on NISQ-Hardware and they additionally required a nested quantum phase estimation for specifically quantum neural networks, we refrained from utilizing this method as a comparison for the studies in our work. Gong et al. [8] proposed a vanishing gradient for introducing adversarial noise on a quantum circuit. Nevertheless, the method of Gong et al. [8] assumes that the data is in a quantum state, which is then adversarially perturbed during the quantum circuit. We focused on encoding classical data on the quantum circuit rather than assuming already encoded quantum states. Hence, in our scenario, the method is hence not applicable. Therefore, we refrained from including Gong et al. [8] in our study. Additionally, Weber et al. [35] have provided a fundamental link between quantum hypothesis testing and binary classification. Their work also shows the connection between classical randomized smoothing and depolarization noise on a quantum circuit. Finally, our approach to train and test the robustness of a multi-class quantum neural network is novel in that the community primarily focuses on binary classification. The only work focusing on the development of variational ansätze for multi-class classifiers is Wang et al. [34], which created a structure to classify MNIST10 with an accuracy of $\approx 90\%$.
@@ 6 Conclusion
Evaluating the robustness of Quantum Neural Networks (QNNs), we addressed three primary research questions.
Firstly, we assessed the impact of depolarization noise on a multi-class classifier. It was observed that introducing more classes to the QNN decreases both training and test accuracy. Furthermore, the overall robustness of the QNN diminishes and is not improved by depolarization noise. This trend is particularly evident when examining the certifiable distances $\tau_{D}$; for all MNIST datasets, $\tau_{D}$ decreases with an increase in depolarizing noise and even tends towards negative values when $p=0.9$ for MNIST4 and MNIST10. Secondly, exploring alternative methods for adversarial training of the network also failed to yield benefits, as did the addition of depolarization noise to adversarially trained networks.
It is important to note that Du et al. [6] asserted that the certifiable distance depends solely on the overall depolarizing noise, the dimensions, and the ratio of the true class probability to the probability of the subsequent class. Our work demonstrates that increasing depolarizing noise in a quantum channel negatively impacts both this certifiable distance and the adversarial accuracy in practice. Consequently, future research should focus on obtaining weights that enhance class separation in the QNN context. This is crucial because neither the number of measurements nor the depolarizing noise in a quantum device can be controlled. This premise initially motivated our exploration of certifiable distance. Hereby it would be interesting to lean on the work of Liao et al. [17], to find more suitable architectures than the one used in this work.
Aside from enhancing quantum neural networks to achieve reasonable performance, a secondary objective could be to replicate this study with other widely-used datasets, such as CIFAR-10 or CIFAR-100. This would be of particular interest, since it could allow a quantum-classical comparison of the robustness of QNNs, towards networks present in Croce et al. [5]  [^7] https://robustbench.github.io/.
Finally, we wish to highlight to the reader that, despite these challenges, the foundational assumption of quantum differential privacy and the feasibility of establishing a robustness bound with minimal depolarization noise in the quantum circuit enable the creation of inherently certifiable networks. However, based on our findings, we acknowledge the limited practical applicability of quantum neural networks in their current form.

### 21 Dec 2023 v2

@@ 4.2 Model Architecture
Table 1: Hyperparameters used for the various networks for the different datasets. We used the qiskit.aer device and pennylanes specs. Single Qubit gates are $R_{z}$ and $R_{y}$, two qubit gates are $CX$.
Hyper-Parameter Iris PID BC MNIST2 MNIST4 MNIST10 CIFAR2 learning rate 0.05 0.0005 0.0005 0.005 0.005 0.005 0.0005 batch size 30 16 16 512 512 128 64 Nr. epochs 100 10 10 30 30 30 10 Nr layers 2 40 40 40 40 40 80 Nr. measurements 1 1 1 1 2 4 1 Nr. Parameters 13 181 481 961 961 961 2641 Circuit Depth 11 121 242 276 276 276 601 Single Qubit Gates 12 180 480 960 960 960 2640 Two Qubit Gates 5 60 160 320 320 320 880
@@
Additionally, due to the noisy architecture and limited availability of quantum devices, we have simulated the quantum device using statevector simulation. Specifically, we utilized the auto-grad feature of pytorch and the pennylane library to formulate the ansatz  [^5] Pennylane [2] 0.29.1 and pytorch [28] 2.0.0.
@@ 5 Experiments
With our experiments, we try to answer the three questions raised above. At first, we want to see the influence of depolarization noise on a multi-class classifier, therefore, we used the Iris, MNIST2, MNIST4 and MNIST10 datasets as well as the PID, BC and CIFAR2 datasets. Secondly, our interest shifts to whether one can defend our QNNs with other adversarial training methods, specifically randomized smoothing and adversarial training (which we refer to as PGD for reasons of conciseness) we hereby focused on the Iris, Breast Cancer, Pima Indians Diabetes, CIFAR2, MNIST2, MNIST4 and MNIST10 datasets. Finally, we want to elaborate on whether we can chain adversarial training with depolarization noise as a blend of two adversarial defense mechanisms, for this particular question we solely evaluated MNIST2.
@@
Dataset Train Param. Acc. Certified Distance
$\tau_{D}$ (×$\times 10^{-2}$)
@@ 5.1 Training of the Models
As already pointed out, our models all follow a AmplitudeEmbedding plus a StronglyEntangling layer structure and only vary in terms of the datasets. We train all models using Adam as an optimizer by varying the batch size, learning rate, and number of epochs. These models are referred to as vanilla models. PyTorch’s categorical-cross entropy was utilized as the loss function for Iris, BC, PID, CIFAR2, MNIST2, and MNIST4. For the MNIST10 dataset, however, the application of negative log likelihood as the loss function, expressed as $l(y,\hat{y})=-\sum_{i}y_{i}\log(\hat{y}_{i})$, was found to enhance accuracy. This effectiveness in the specific case of MNIST10 led to its selection as the preferred loss function for this dataset. Regarding adversarial training, we train the models using the PGD attack with $L_{\infty}$-norm for the Iris dataset with $T=50$ steps. Due to performance issues, we train the MNIST2, MNIST4, and MNIST10 models with $T=3$ steps. We also employ randomized smoothing during training; thus, we perturb the original image with noise from a uniform distribution and use this as a training sample. The variance of this training is reported as $\sigma$, and the mean is set to zero. We employ early stopping for both randomized smoothing and adversarial training. We use the loss value at which the models are considered to be converged. For Iris, this value is set to $0.6$. For MNIST2, the value is set to $0.45$, for MNIST4 to $1.15$, and for MNIST10 to $1.7$. Note that we do not perform adversarial training on BC, PID, or CIFAR2, since these datasets are solely used to assess performance under depolarization noise. The test accuracy, combined with the mean certified distance $\tau_{D}$ derived from the test set, is reported in Table 2.
@@
(b) BC
(c) PID
(d) MNIST2
(e) MNIST4
(f) MNIST10
Figure 4: Robustness accuracy for the various vanilla classifiers for Iris, BC, PID, MNIST2, MNIST4 and MNIST10.
5.2 Effect of Depolarization Noise
We compare the vanilla networks of the Iris, PID, BC, MNIST2, MNIST4, and MNIST10 datasets for depolarization noises $p$ in the range $[0.1,0.5,0.9]$. We include a depolarizing noise of $0.3$ for MNIST2. Our results are shown in Figure 4. Additionally, we report the adversarial accuracy of CIFAR2 for the depolarization factors of $[0.1,0.5,0.9]$ for $\epsilon\in[0.0,0.3]$. We choose the factor of $0.3\approx\nicefrac{{8}}{{255}}$, as this is also the factor used in the benchmarks conducted on CIFAR10. However, we note that our version of CIFAR2 is not comparable with the benchmarks done on CIFAR10, since we have converted the multi-class classification into a binary classification and our attack is conducted on the HOG features rather than on the actual pixels.
Regarding the binary classification networks BC, PID, Iris, and MNIST2, we typically observe that the QNN performs well in terms of accuracy. With PID, this is noteworthy as the QNN achieves only an accuracy of $67.08\%$. Comparing this score to the random forest approach in Chang et al. [7], which resulted in an accuracy of $79.57\%$, we see that this accuracy corresponds to the upper quartile of the experiments, as shown in Egginger et al. [10, Fig. 1].
For Iris, BC, and PID, we observe a similar behavior as seen in Du et al. [9, Apdx. H]: Additional depolarization noise typically aids the classifier for larger epsilon values, although too much noise might decrease the initial accuracy. PID is particularly interesting, as it shows this effect even for small adversarial perturbations. However, the practical use of this effect is questionable, as it typically occurs when the classifier is already below or close to $50\%$ accuracy.
In contrast, for MNIST2, we do not find any significant improvement in terms of adversarial accuracy. A similar behavior is observed for CIFAR2, as detailed in table Table 3. For MNIST4 and MNIST10, we note that depolarization noise does not increase robustness. This is attributed to two factors.
First, the limited capacity of the network to learn the data distribution. This is supported by Table 2, showing that accuracy decreases with an increase in the number of classes. Additionally, when examining the certified bounds, we notice that the minimum of all computed bounds is typically negative, and the maximum trace distance is usually smaller than for binary classification. The negative bound is attributed to our inability to achieve $100\%$ accuracy on the test set.
Secondly, for MNIST4 and MNIST10, we measure two qubits to obtain the classes for MNIST4 and four qubits for MNIST10. The poor performance could also be attributed to depolarization noise automatically increasing the probability of the follow-up class, since two classes are assigned to one qubit. However, this assumption requires further experimental and analytical evaluation, which is a subject for future work.
Table 3: Mean Robustness Accuracy for CIFAR2 for various depolarization noises
Depol. Robustness Accuracy (%)
$p$ $\epsilon=0.0$ $\epsilon=0.03$
0.0 82.5 56.0
0.1 79.0 58.0
0.5 49.0 49.0
0.9 49.0 49.0
@@ 5.3 Effect of Adversarial Training
Figure 5: Comparison of the two methods randomized smoothing (RS) and adversarial training (PGD) specifically for the Iris dataset.
Here, we observe the effect of adversarial training and randomized smoothing on the Iris dataset (see Figure 5). Overall, we find that adversarial training itself has no influence for the selected epsilon values. In contrast, for a large radius in randomized smoothing, the adversarial accuracy drops faster initially but remains more resilient for larger epsilon values. As a next step, we conduct the same experiments on the MNIST2, MNIST4, and MNIST10 datasets, to examine the influence of adversarial training and randomized smoothing for a larger set of classes. Our results are visualized in Figure 6. We observe that neither adversarial training nor randomized smoothing significantly enhances the robustness of the QNN. Additionally, a similar behavior to that seen with depolarizing noise is evident, specifically that the networks tend to become less robust with larger classes. This trend parallels the fact that the testing accuracy drops by approximately 10% from MNIST2 to MNIST4 and around 10% from MNIST4 to MNIST10. Interestingly, randomized smoothing and adversarial training do not tend to worsen the robustness accuracy.
5.4 Effect of Depolarization Noise and Adversarial Training
Since adversarial training and the effect of depolarization noise behave similarly, we explore whether adding depolarization noise on top of an adversarial training procedure can improve the general robustness of the network. Therefore, we closely examine the MNIST2 dataset and enhance the adversarially trained QNN circuits with depolarization noise while attacking them. Our results are observed in Figure 7. Overall, we find that adding depolarization noise on top of adversarial training does not provide improvements for the MNIST2 dataset. A similar picture emerges for other datasets when examining the certified trace distances on the test set in Table 2.
(a) MNIST2: PGD
(b) MNIST4: PGD
(c) MNIST10: PGD
@@ 5.4 Effect of Depolarization Noise and Adversarial Training
(e) MNIST4: RS
(f) MNIST10: RS
Figure 6: Comparison of the two methods randomized smoothing (RS) and adversarial training (PGD) for a set of parameters on the QNN. Adversarial training was conducted with PGD($L_{\infty}$, $T=50$) for the Iris classifier, PGD($L_{\infty}$, $T=3$) for all MNIST classifiers
@@ 6 Conclusion
Evaluating the robustness of QNNs, we addressed three primary research questions. We assessed the impact of depolarization noise on a quantum classifier. For the binary classifiers, we were able to observe that adding depolarizing noise might improve the robustness accuracy, nevertheless only for larger epsilon values and typically after the classifier was perturbed below the $50\%$ mark. We doubt that this behavior is helpful in practice, since one would like to provide a classifier that is more robust for smaller epsilon values. Next, it was observed that introducing more classes to the QNN decreases both training and test accuracy. Furthermore, the overall robustness of the QNN diminishes and is not improved by depolarization noise. This trend is particularly evident when examining the certifiable distances $\tau_{D}$; for all MNIST datasets, $\tau_{D}$ decreases with an increase in depolarizing noise and even tends towards negative values when $p=0.9$ for MNIST4 and MNIST10.
Exploring alternative methods for adversarial training of the network also failed to yield benefits, as did the addition of depolarization noise to adversarially trained networks.
It is important to note that Du et al. [9] asserted that the certifiable distance depends solely on the overall depolarizing noise, the dimensions, and the ratio of the true class probability to the probability of the subsequent class. Our work demonstrates that increasing depolarizing noise in a quantum channel negatively impacts both this certifiable distance and the adversarial accuracy in practice. Consequently, future research should focus on obtaining weights that enhance class separation in the QNN context. This is crucial because neither the number of measurements nor the depolarizing noise in a quantum device can be controlled. This premise initially motivated our exploration of certifiable distance. In this regard, it would be interesting to build on the work of Liao et al. [21], to find more suitable architectures than the one used in this work.